OpenStudy (zubhanwc3):

write a function so that as the limit of x approaches 0 of f(x) it doesnt exist, but as the limit of x approaches 0, of the absolute value of f(x) exists

OpenStudy (anonymous):

make a piecewise function might work

OpenStudy (zubhanwc3):

write a function so that \[\lim_{x \rightarrow 0}f(x)\] doesnt exist, but \[\lim_{x \rightarrow 0} \left| f(x) \right|\] exists

OpenStudy (zubhanwc3):

thats wat i thought, but i cant think of one as x approaches 0, maybe for 2 or 3, but not 0 :(

OpenStudy (anonymous):

you want a simple answer, or an idea?

OpenStudy (zubhanwc3):

an answer, and explanation maybe? ive been stuck on this problem for 3 days :(

OpenStudy (anonymous):

ok first of all, all functions you know are continuous on their domain all trig functions, exponential functions, rational functions logarithmic function etc etc

OpenStudy (anonymous):

so if you want to come up with an example of a function where the limit at 0 or any other number does not exist, your best bet is to come up with a piecewise function, defined one way for \(x<0\) and another way for \(x>0\)

OpenStudy (anonymous):

you know what i mean when i say a "piecewise" function?

OpenStudy (anonymous):

"no" is a fine answer, i am just asking i can explain if you like

OpenStudy (zubhanwc3):

yes i do, its usually used with absolute value functions, or filling in a removable discontinuity

OpenStudy (anonymous):

here is an example \[f(x) = \left\{\begin{array}{rcc} x + 4 & \text{if} & x <0 \\ - x - 4& \text{if} & x\geq 0 \end{array} \right. \]

OpenStudy (zubhanwc3):

wouldnt the top one be > rather than <

OpenStudy (anonymous):

as a matter of fact, that might be an example you could use as an answer i was going to give a much simpler one so simple you are going to be annoyed that you spent three days or even three minutes on it

OpenStudy (anonymous):

no they can't both be greater than

OpenStudy (zubhanwc3):

no, i mean the bottom one would be < and equal to, while the top one is >

OpenStudy (anonymous):

how about this simple function \[f(x) = \left\{\begin{array}{rcc} -1& \text{if} & x <0 \\ 1& \text{if} & x\geq 0 \end{array} \right. \]

OpenStudy (anonymous):

it is my function i can write it any way i choose

OpenStudy (zubhanwc3):

........... but then the function wouldnt be continuous with that function

OpenStudy (zubhanwc3):

i mean, limit wouldnt exist

OpenStudy (anonymous):

write a function so that as the limit of x approaches 0 of f(x) it doesnt exist,

OpenStudy (anonymous):

if that is the case, it cannot be continuous

OpenStudy (zubhanwc3):

because the left hand limit wouldn't equal the right hand limit

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} -1& \text{if} & x <0 \\ 1& \text{if} & x\geq 0 \end{array} \right.\] the left handed limit is \(-1\) and the right handed limit is \(1\)

OpenStudy (zubhanwc3):

and one doesn't equal -1, also, there needs to be a non absolute value function where the limit wouldnt exist

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

clear or no? first it says \(\textbf{"write a function so that as the limit of x approaches 0 of f(x) it doesnt exist"}\) i say \[f(x) = \left\{\begin{array}{rcc} -1& \text{if} & x <0 \\ 1& \text{if} & x\geq 0 \end{array} \right.\] fits that bill

OpenStudy (zubhanwc3):

o wait, i get it now

OpenStudy (zubhanwc3):

by being absolute value, both sides would be 1 rather than -1 and 1 o.o

OpenStudy (anonymous):

then it says but as the limit of x approaches 0, of the absolute value of f(x) exists i say \[f(x) = \left\{\begin{array}{rcc} -1& \text{if} & x <0 \\ 1& \text{if} & x\geq 0 \end{array} \right.\] als fits that bill because \[|f(x)|=1\]

OpenStudy (zubhanwc3):

thank you, u were right, a simple function solved 3 days of effort T_T

OpenStudy (anonymous):

make up your own example now

OpenStudy (zubhanwc3):

ok

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} x + 4 & \text{if} & x <0 \\ - x - 4& \text{if} & x\geq 0 \end{array} \right.\] might also work, you can check it